OpenStudy (anonymous):

the owners of Nuts2U snack shack mixes cashews worth 6 dollars a pound with peanuts worth $2.10 a pound to get a half-pound mixed nut bag worth $1.70. how much of watch kind of nut is included in the mixed bag?

OpenStudy (studygurl14):

Okay, so use these equations: c + p = 0.5 lb 6c + 2.10p = 1.70 (c = number of cashews p = number of peanuts )

OpenStudy (studygurl14):

plug in the info you know. and solve

OpenStudy (studygurl14):

c = 0.5 - p p = 0.5 -c Plug in

OpenStudy (anonymous):

studygur you know how to solve this 4x-5<-9or3x+4>7

OpenStudy (studygurl14):

yes, I sorta remember how to do it. but I have to leave soon.

OpenStudy (studygurl14):

I'll start it thogh with you

OpenStudy (anonymous):

thanks study girl!:))

OpenStudy (studygurl14):

you're welcome @calebthorne

OpenStudy (anonymous):

i put in the price of them per pound? @StudyGurl14

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

@calebthorne do this 6c + 2.10p = 1.70 6(0.5 - p) + 2.10p = 1.70 You see?

OpenStudy (studygurl14):

@mayabdo you with me?

OpenStudy (studygurl14):

I'll start yours

OpenStudy (studygurl14):

First, rearrange the first one: 4x-5<-9 Isolate the x 4x< -4 What next to isolate x?

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

What next o isolate the x, @mayabdo ?

OpenStudy (anonymous):

i think that 4x-4=

OpenStudy (studygurl14):

no, it is an inequality

OpenStudy (studygurl14):

divide both sides by 4 and get...?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

than

OpenStudy (anonymous):

you with me

OpenStudy (studygurl14):

x < -1